OpenStudy (anonymous):

Perform the indicated oeration and write in standard form: (4+√-2)(5+√-5) Thats a square root sign

OpenStudy (anonymous):

\[=\frac{4+i \sqrt{2}}{5+i \sqrt{5}}*\frac{5-i \sqrt{5}}{5-i \sqrt{5}}\]now multiply numerator and denomnator (denominator is a sum of squares) Continued on next post...

OpenStudy (anonymous):

FOIL in numerator; sum of squares in denominator\[=\frac{20-4i \sqrt{5}+5i \sqrt{2}+ \sqrt{10}}{25+5}\]next simplify numerator and denominator\[=\frac{(20+\sqrt{10})+(5\sqrt{2}-4\sqrt{5})i}{30}\] and you are done

OpenStudy (anonymous):

You could write the real and imaginary parts each over the 30 and that would be standard form, which some teachers/books want you to do.